### v1

@@ 1 Introduction
Valiant [27] proposed an algebraic version of the $\mathsf{P}$ versus $\mathsf{NP}$ question and defined the class $\mathsf{VP}$, the algebraic analogue of $\mathsf{P}$, which contains polynomial families computable by polynomial sized algebraic circuits. An algebraic circuit (or, arithmetic circuit) $C$ is a directed acyclic graph such that (1) every node has either in-degree (fan-in) 0 (the input gates) or $2$ (the computational gates), (2) every input gate is labeled by elements from a field $\mathbb{K}$ or variables from $\mathbf{X}=\{X_{1},\cdots,X_{n}\}$, (3) every computational gate is labeled by either ++ (addition gate) or $\times$ (multiplication gate), with the obvious syntactic meaning, and (4) there is a unique gate of out-degree 0, the output gate. Clearly, every gate in a circuit computes a polynomial in $\mathbb{K}[\mathbf{X}]$. We say that the circuit $C$ computes $P(\mathbf{X})\in\mathbb{K}[\mathbf{X}]$ if the output gate of $C$ computes $P(\mathbf{X})$. The size of $C$, denoted by ${\rm size}(C)$, is the number of nodes in the circuit. An algebraic circuit is an algebraic formula if every gate in the circuit has out-degree $1$ except for the output gate. The class $\mathsf{VNP}$, the algebraic analogue of $\mathsf{NP}$, is definable by taking exponential sums of the form
@@
where $g$ is computable by a polynomial-size circuit and $\ell$ is polynomial in the number of variables. It is known that one can also replace algebraic circuits by algebraic formulas, and still get the same class $\mathsf{VNP}$ [27, 22]. Valiant further proved that the permanent family is complete for $\mathsf{VNP}$ (over fields of characteristic not two). Recall that the permanent of a matrix $(X_{i,j})$ is defined as
@@
The famous Valiant’s conjecture $\mathsf{VP}\not=\mathsf{VNP}$ is equivalent to the fact that the permanent does not have polynomial-size circuits. The representation of the permanent in (2), although it looks very natural, is not optimal. Ryser’s formula [23] yields an algebraic formula of size $O(2^{n}n^{2})$. A formula of similar size was later found by Glynn [15]. Ryser’s formula is now over sixty years old and has not been improved since. This gives rise to the interesting question whether there is a formula or circuit of subexponential-size (in $n$) for the permanent? More generally, we can now ask the following question.
@@
Note that exponential-size being necessary is a much stronger claim than $\mathsf{VP}\not=\mathsf{VNP}$. It could well be that $\mathsf{VP}\not=\mathsf{VNP}$ but still exponential sums like in (1) have subxponential size circuits! In this paper, we shed some light on the question what happens if exponential sums would always have subexponential size circuits.
Question 1.1 works as driving force between the famous Shub-Smale $\tau$-conjecture [24] and exponential lower bounds on exponential sums. The $\tau$-complexity $\tau(f)$ of an integer polynomial is the size of a smallest division-free circuit that computes $f$ starting from the constants $\pm 1$. The $\tau$-conjecture states the the number of integer zeroes of $f$ is polynomially bounded in $\tau(f)$, see [24]. [24] shows that the $\tau$-conjecture implies $\mathsf{P}_{\mathbb{C}}\neq\mathsf{NP}_{\mathbb{C}}$, in the Blum–Shub–Smale (BSS) model of computation over the complex numbers [6, 5].
Super-polynomial lower bounds assuming the $\tau$-conjecture. Bürgisser [9] connected the $\tau$-complexity of the permanent to various other conjectures. He showed that the $\tau$-conjecture implies a superpolynomial lower bound on $\tau(\operatorname{per}_{n})$, implying the constant-free version of $\mathsf{VP}\neq\mathsf{VNP}$, namely $\mathsf{VP}^{0}\neq\mathsf{VNP}^{0}$; for definitions, see Section˜2.1. The proof strategy of [9] is as follows: assume $\tau(\operatorname{per}_{n})={\rm poly}(n)$, and conclude a complexity-theoretic ‘collapse’ that the counting hierarchy $\mathsf{CH}$ (for a definition, see Section˜3) is in $\mathsf{P}/{\rm poly}$. Consider the Pochhammer–Wilkinson polynomial $f_{n}(x):=\prod_{i=1}^{n}(x-i)$, and construct a unique $O(\log n)$-variate multilinear polynomial $B_{n}$ such that under a ‘suitable’ substitution, one gets back $f_{n}$. The coefficients of $f_{n}$ as well as $B_{n}$, are efficiently computable (since $\mathsf{CH}\subseteq\mathsf{P}/{\rm poly}$), implying $B_{n}\in\mathsf{VNP}^{0}$. An inspection of Valiant’s completeness result reveals that if $B_{n}\in\mathsf{VNP}^{0}$, then there is a polynomially bounded sequence $p(n)$ such that $\tau(2^{p(n)}B_{n})={\rm poly}(\log n)$, which implies $\tau(2^{p(n)}f_{n})={\rm poly}(\log n)$, contradicting the $\tau$-conjecture.
In [9], the superpolynomial lower bound on $\tau(\operatorname{per}_{n})$ was also implied by any of the quantities $\tau(n!)$, $\tau(\sum_{k=0}^{n}\frac{1}{k!}T^{k})$, or $\tau(\sum_{k=0}^{n}k^{r}T^{k})$ (for any fixed negative integer $r$) not being poly-logarithmically bounded as a function of $n$. Here, we remark that the separation proof of $\mathsf{VP}^{0}$ and $\mathsf{VNP}^{0}$, even assuming strong bounds on the $\tau$-conjecture, is merely superpolynomial: we do not get the (possibly) desirable exponential separation between $\mathsf{VP}^{0}$ and $\mathsf{VNP}^{0}$. This leads to the following question.
@@
Here, we mention that there are variants of the $\tau$-conjecture, e.g., the real $\tau$-conjecture [19, 25] or SOS-$\tau$-conjecture [12], which also give strong algebraic lower bounds. There is also super polynomial lower bound known from a proof complexity theoretic view due to [1] from the original Shub-Smale $\tau$-conjecture. However, the Shub-Smale $\tau$-conjecture is not known to give an exponential lower bound for the permanent.
@@ 1.1 Our results
The results of our paper revolve around answering both ˜1.1 and 1.2 positively. The main result is the following.
Theorem 1.3 (Informal).
@@
(3) Since there’s a polynomial (non-linear) blowup in the reduction of the exponential sum on the universal circuit from the permanent, we will only get a subexponential lower bound on the permanent polynomial assuming the $\tau$-conjecture. We leave it as an open question to achieve an exponential lower bound on the permanent assuming the $\tau$-conjecture.
@@
where $k=n/\log m$. The size of the exponential is measured in the number $\ell(n)$ of $\mathbf{Y}$-variables. In the end, we want to measure in the input size, the number $n$ of $\mathbf{X}$-variables. To talk about subexponential complexity, $\ell(n)$ should be linearly bounded. $g$ will be typically computed by a circuit (of unbounded degree). We want to view $\mathsf{p\text{-}{{\log}\text{-}Expsum}}_{m,k}$ as a parameterized problem, the parameter will be $k$. Our definition of $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$, as a polynomial-sum, is motivated by the log-parameterizations which are used in the definition of the so-called $M$-hierarchy in the Boolean setting, see [13, 14].
We show that $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is most likely not fixed-parameter tractable (fpt). A polynomial family $p_{n,k}$ is fpt if both its size and degree are fpt bounded, i.e., of the form $f(k)q(n)$, for $q\leq{\rm poly}(n)$, and $f:\mathbb{N}\to\mathbb{N}$ being any computable function. We connect $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ with – (1) a linear variant of the counting hierarchy (we denote it by $\mathsf{CH}_{\mathrm{lin}}$), where the size of the oracle calls are bounded linearly in the size of the input; for definition see Section 3; and (2) integers definable in $\mathsf{CH}_{\mathrm{lin}}$, similar to Bürgisser [9]. Informally, an integer is definable in $\mathsf{CH}_{\mathrm{lin}}$, if its sign and bits are computable in the same class.
Theorem 1.4 (Informal).
@@
For formal statements, see Theorem 4.1 and 6.1.
Finally, many algebraic complexity classes can be defined in terms of permanents. Most prominently, the “regular” permanent family $(\operatorname{per}_{n})$ is complete for $\mathsf{VNP}$. The class $\mathsf{VW}[1]$ is an important class in parameterized complexity. It is defined as a bounded sum over constant depth weft-1 circuits. Bounded sum means that we sum over $\{0,1\}$-vectors with $k$ ones and $k$ is the parameter. Bläser and Engels [4] prove that $\mathsf{VW}[1]$ is described by so-called $k$-permanents with $k$ being the parameter. In a $k$-permanent, we only sum over permutations with nkn-k self-loops. The crucial parameterized class of this work is $\mathsf{VW}[\mathsf{P}]$: it is defined as a bounded exponential sum over polynomially-sized arithmetic circuits computing a polynomial of degree that is polynomially bounded. While we do not characterise $\mathsf{VW}[\mathsf{P}]$ in terms of permanents, we characterize the related class $\mathsf{VW}[\mathsf{F}]$: Here instead of summing over circuits, we sum over formulas. [^1] Maybe an explanation of the naming convention is helpful: In $\mathsf{VW}[\mathsf{P}]$, we sum of polynomial-size circuits, which describe the class $\mathsf{VP}$. In $\mathsf{VW}[\mathsf{F}]$, we sum over polynomial size formulas, which define the class $\rm VF$}, the modern name for $\mathsf{VP}_{e}$. The permutations that we sum over for defining our permanent family will have one cycle of length $k$ and all other cycles bounded by $4$. Again, $k$ is the parameter. We call the corresponding polynomials $(k,4)$-restricted permanents. It turns out that we also need to restrict the graph classes. We call a graph $G=(V,E)$ $(4,b)$-nice if we can partition the set $V=V_{1}\cup V_{2}$ disjointly, such that in the induced graph $G[V_{1}]$, every cycle is either a self-loop or has length >4>4 and in the induced graph $G[V_{2}]$ has tree-width bounded by $b$. While this looks artificial at a first glance, it turns out that there is a constant $b$ such that $(k,4)$-restricted permanent on $(4,b)$-nice graphs describes the natural class $\mathsf{VW}[\mathsf{F}]$. There is a family of $(4,b)$-nice graphs such that the corresponding family of $(k,4)$-restricted permanents is $\mathsf{VW}[\mathsf{F}]$-hard. On the other hand, the $(k,4)$-restricted permanent family is in $\mathsf{VW}[\mathsf{F}]$ for every family of $(4,b)$-nice graphs. Together, this implies:
@@ 1.2 Proof ideas
Proof of Part (1): We prove even a stronger statement for the subexponential version of the linear counting hierarchy. The proof goes via induction on the level of the counting hierarchy. The criteria for some language $B$ being in the $(k+1)$-th level is that there should be some language $A$ in the $k$-th level such that $|\{y\in\{0,1\}^{n}:\langle x,y\rangle\in A\}|>2^{n-1}$. Essentially, for a language $A$ in the $k$-th level, we express $|\{y\in\{0,1\}^{n}:\langle x,y\rangle\in A\}|>2^{n-1}$ as an exponential sum over an algebraic circuit $\chi_{A}(x,y)$, which captures the characteristic function of $A$. Furthermore, one can show that $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is fpt (in an unbounded constant-free setting) iff $\sum_{y}g(\mathbf{X},y)$ has $2^{o(n)}{\rm poly}(m)$ size circuits, where $g$ has a circuit of size $m$; see Theorem˜4.3 and 4.4. Putting these together, one gets that the exponential sum has a subexponential-size constant-free circuit. Lastly, we want to get the information about the highest bit of the sum (which is equivalent to looking at it $\bmod\penalty 10000\ 2^{n}$), which can be efficiently arithmetized. In every step there is polynomial blowup in the size, and hence the size remains subexponential, yielding the desired result. For details, see Theorem˜4.1.
@@
Proof of Part (2): This proof is an adaption of [9, 18] in our context. Take a sequence $(a_{n})_{n}\in\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. We define a multilinear polynomial $A(\mathbf{Y})$ such that the coefficient of $\mathbf{Y}^{\mathbf{j}}$ is the $j$-th bit of $a(n)$, where $\mathbf{j}$ is the binary representation of $j$. Furthermore, checking $a(n,j)=b$ can be done by a subexponential circuit $C(\mathbf{N},\mathbf{J})$, where $\mathbf{N}$ and $\mathbf{J}$ have $\log n$ and $\mathsf{bit}(n)$-many variables capturing $n$ and $j$ respectively. Moreover, one can define $F(\mathbf{N},\mathbf{Y},\mathbf{J})=C(\mathbf{N},\mathbf{J})\cdot\prod_{i}(J_{i}Y_{i}+1-J_{i})$ and show that $A$ can be expressed as an exponential sum over $F(j,\mathbf{N},\mathbf{Y})$! This is clearly a $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ instance, which finally yields that the $\tau$-complexity of $a(n)$ is subpolynomial. A similar proof strategy also holds for the polynomials with coefficients being definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. For details, see Section˜6.
@@
Proof idea of Theorem˜1.3 Take the Pochhammer polynomial $p_{n}(X)=\prod_{i=1}^{n}(X+i)$. The coefficient of $X^{n-k}$ in $p_{n}$ will be $\sigma_{k}(1,\dots,n)$, where $\sigma_{k}(z_{1},\dots,z_{n})$ is the $k$-th elementary symmetric polynomial in variables $z_{1},\dots,z_{n}$. It is not hard to show that $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$ is closed under polynomially-many additions and multiplications (Theorem˜5.3). Therefore, $(\sigma_{k}(1,\dots,n))_{n\in\mathbb{N},k\leq n}$ is definable in the linear counting hierarchy (see Corollary˜5.4). And by Theorem˜6.1, $(p_{n})_{n\in\mathbb{N}}$ has $n^{o(1)}$-sized constant-free circuits if $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is fixed-parameter tractable. But $p_{n}$ has $n$ distinct integer roots. Assuming the $\tau$-conjecture, $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is not fpt. On the other hand, one can show that when exponential sums over circuits of size $m$ have circuits have size $2^{o(n)}{\rm poly}(m)$, then the $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is fpt, by Theorem 4.4; in other words, $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is not fpt implies an exponential lower bound on an exponential sum. This finishes the proof.

### v2

@@ 1 Introduction
Valiant [VAL79] proposed an algebraic version of the $\mathsf{P}$ versus $\mathsf{NP}$ question and defined the class $\mathsf{VP}$, the algebraic analogue of $\mathsf{P}$, which contains polynomial families computable by polynomial sized algebraic circuits. An algebraic circuit (or, arithmetic circuit) $C$ is a directed acyclic graph such that
  1. 1.
    every node has either in-degree (fan-in) 0 (the input gates) or $2$ (the computational gates),
  2. 2.
    every input gate is labeled by elements from a field $\mathbb{K}$ or variables from $\mathbf{X}=\{X_{1},\cdots,X_{n}\}$,
  3. 3.
    every computational gate is labeled by either ++ (addition gate) or $\times$ (multiplication gate), with the obvious syntactic meaning, and
  4. 4.
    there is a unique gate of out-degree 0, the output gate.
Clearly, every gate in a circuit computes a polynomial in $\mathbb{K}[\mathbf{X}]$. We say that the circuit $C$ computes $P(\mathbf{X})\in\mathbb{K}[\mathbf{X}]$ if the output gate of $C$ computes $P(\mathbf{X})$. The size of $C$, denoted by ${\rm size}(C)$, is the number of nodes in the circuit. An algebraic circuit is an algebraic formula if every gate in the circuit has out-degree $1$ except for the output gate. The class $\mathsf{VNP}$, the algebraic analogue of $\mathsf{NP}$, is definable by taking exponential sums of the form
@@
where $g$ is computable by a polynomial-size circuit and $\ell$ is polynomial in the number of variables. It is known that one can also replace algebraic circuits by algebraic formulas, and still get the same class $\mathsf{VNP}$ [VAL79, MP08]. Valiant further proved that the permanent family is complete for $\mathsf{VNP}$ (over fields of characteristic not two). Recall that the permanent of a matrix $(X_{i,j})$ is defined as
@@
The famous Valiant’s conjecture $\mathsf{VP}\not=\mathsf{VNP}$ is equivalent to the fact that the permanent does not have polynomial-size circuits. The representation of the permanent in (2), although it looks very natural, is not optimal. Ryser’s formula [RYS63] yields an algebraic formula of size $O(2^{n}n^{2})$. A formula of similar size was later found by Glynn [GLY10]. Ryser’s formula is now over sixty years old and has not been improved since. This gives rise to the interesting question whether there is a formula or circuit of subexponential-size (in $n$) for the permanent? More generally, we can now ask the following question.
@@
Note that exponential-size being necessary is a much stronger claim than $\mathsf{VP}\not=\mathsf{VNP}$. It could well be that $\mathsf{VP}\not=\mathsf{VNP}$ but still exponential sums like in (1) have subexponential size circuits! In this paper, we shed some light on the question what happens if exponential sums would always have subexponential size circuits.
Question 1.1 serves as a driving force between the famous Shub-Smale $\tau$-conjecture [SS95] and exponential lower bounds on exponential sums. The $\tau$-complexity $\tau(f)$ of an integer polynomial is the size of a smallest division-free circuit that computes $f$ starting from the constants $\pm 1$. The $\tau$-conjecture states that the number of integer zeroes of $f$ is polynomially bounded in $\tau(f)$, see [SS95]. [SS95] shows that the $\tau$-conjecture implies $\mathsf{P}_{\mathbb{C}}\neq\mathsf{NP}_{\mathbb{C}}$, in the Blum–Shub–Smale (BSS) model of computation over the complex numbers [BSS89, BCS+00].
Super-polynomial lower bounds assuming the $\tau$-conjecture. Bürgisser [BÜR09] connected the $\tau$-complexity of the permanent to various other conjectures. He showed that the $\tau$-conjecture implies a superpolynomial lower bound on $\tau(\operatorname{per}_{n})$, implying the constant-free version of $\mathsf{VP}\neq\mathsf{VNP}$, namely $\mathsf{VP}^{0}\neq\mathsf{VNP}^{0}$; for definitions, see Section˜2.1. The proof strategy of [BÜR09] is as follows: assume $\tau(\operatorname{per}_{n})={\rm poly}(n)$, and conclude a complexity-theoretic ‘collapse’ that the counting hierarchy $\mathsf{CH}$ (for a definition, see Section˜3) is in $\mathsf{P}/{\rm poly}$. Consider the Pochhammer–Wilkinson polynomial $f_{n}(x):=\prod_{i=1}^{n}(x-i)$, and construct a unique $O(\log n)$-variate multilinear polynomial $B_{n}$ such that under a ‘suitable’ substitution, one gets back $f_{n}$. The coefficients of $f_{n}$ as well as $B_{n}$, are efficiently computable (since $\mathsf{CH}\subseteq\mathsf{P}/{\rm poly}$), implying $B_{n}\in\mathsf{VNP}^{0}$. An inspection of Valiant’s completeness result reveals that if $B_{n}\in\mathsf{VNP}^{0}$, then there is a polynomially bounded sequence $p(n)$ such that $\tau(2^{p(n)}B_{n})={\rm poly}(\log n)$, which implies $\tau(2^{p(n)}f_{n})={\rm poly}(\log n)$, contradicting the $\tau$-conjecture.
In [BÜR09], the superpolynomial lower bound on $\tau(\operatorname{per}_{n})$ was also implied by any of the quantities $\tau(n!)$, $\tau(\sum_{k=0}^{n}\frac{1}{k!}T^{k})$, or $\tau(\sum_{k=0}^{n}k^{r}T^{k})$ (for any fixed negative integer $r$) not being poly-logarithmically bounded as a function of $n$. Here, we remark that the separation proof of $\mathsf{VP}^{0}$ and $\mathsf{VNP}^{0}$, even assuming strong bounds on the $\tau$-conjecture, is merely superpolynomial: we do not get the (possibly) desirable exponential separation between $\mathsf{VP}^{0}$ and $\mathsf{VNP}^{0}$. This leads to the following question.
@@
Here, we mention that there are variants of the $\tau$-conjecture, e.g., the real $\tau$-conjecture [KOI11, TAV14] or SOS-$\tau$-conjecture [DUT21], which also give strong algebraic lower bounds. There is also super polynomial lower bound known from a proof complexity theoretic view due to [AGH+19] from the original Shub-Smale $\tau$-conjecture. However, the Shub-Smale $\tau$-conjecture is not known to give an exponential lower bound for the permanent.
@@ 1.1 Our results
The results of our paper revolve around answering both ˜1.1-1.2 positively. The main result is the following.
Theorem 1.3 (Informal version of Theorem˜6.4).
@@
(3) Since there is a polynomial (non-linear) blowup in the reduction of the exponential sum on the universal circuit from the permanent, we will only get a subexponential lower bound on the permanent polynomial assuming the $\tau$-conjecture. We leave it as an open question to achieve an exponential lower bound on the permanent assuming the $\tau$-conjecture.
@@
where $k=n/\log m$. The size of the exponential is measured in the number $\ell(n)$ of $\mathbf{Y}$-variables. In the end, we want to measure in the input size, the number $n$ of $\mathbf{X}$-variables. To talk about subexponential complexity, $\ell(n)$ should be linearly bounded. $g$ will be typically computed by a circuit (of unbounded degree). We want to view $\mathsf{p\text{-}{{\log}\text{-}Expsum}}_{m,k}$ as a parameterized problem, the parameter will be $k$. Our definition of $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$, as a polynomial-sum, is motivated by the log-parameterizations which are used in the definition of the so-called $M$-hierarchy in the Boolean setting, see [FG04, FG06].
We show that $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is most likely not fixed-parameter tractable (fpt). A polynomial family $p_{n,k}$ is fpt if both its size and degree are fpt bounded, i.e., of the form $f(k)q(n)$, for $q\leq{\rm poly}(n)$, and $f:\mathbb{N}\to\mathbb{N}$ being any computable function. We connect $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ with
  1. 1.
    a linear variant of the counting hierarchy (we denote it by $\mathsf{CH}_{\mathrm{lin}}$), where the size of the oracle calls are bounded linearly in the size of the input; for definition see Section 3; and
  2. 2.
    integers definable in $\mathsf{CH}_{\mathrm{lin}}$, similar to Bürgisser [BÜR09]. Informally, an integer is definable in $\mathsf{CH}_{\mathrm{lin}}$, if its sign and bits are computable in the same class.
Theorem 1.4 (Informal version of Theorem˜4.1 and Theorem˜6.1).
@@
Finally, many algebraic complexity classes can be defined in terms of permanents. Most prominently, the “regular” permanent family $(\operatorname{per}_{n})$ is complete for $\mathsf{VNP}$. The class $\mathsf{VW}[1]$ is an important class in parameterized complexity. It is defined as a bounded sum over constant depth weft-1 circuits. Bounded sum means that we sum over $\{0,1\}$-vectors with $k$ ones and $k$ is the parameter. Bläser and Engels [BE19] prove that $\mathsf{VW}[1]$ is described by so-called $k$-permanents with $k$ being the parameter. In a $k$-permanent, we only sum over permutations with nkn-k self-loops.
The crucial parameterized class of this work is $\mathsf{VW}[\mathsf{P}]$: it is defined as a bounded exponential sum over polynomially-sized arithmetic circuits computing a polynomial of degree that is polynomially bounded. While we do not characterise $\mathsf{VW}[\mathsf{P}]$ in terms of permanents, we characterize the related class $\mathsf{VW}[\mathsf{F}]$: Here instead of summing over circuits, we sum over formulas. [^1] Maybe an explanation of the naming convention is helpful: In $\mathsf{VW}[\mathsf{P}]$, we sum of polynomial-size circuits, which describe the class $\mathsf{VP}$. In $\mathsf{VW}[\mathsf{F}]$, we sum over polynomial size formulas, which define the class $\rm VF$}, the modern name for $\mathsf{VP}_{e}$. The permutations that we sum over for defining our permanent family will have one cycle of length $k$ and all other cycles bounded by $4$. Again, $k$ is the parameter. We call the corresponding polynomials $(k,4)$-restricted permanents.
It turns out that we also need to restrict the graph classes. We call a graph $G=(V,E)$ $(4,b)$-nice if we can partition the set $V=V_{1}\cup V_{2}$ disjointly, such that in the induced graph $G[V_{1}]$, every cycle is either a self-loop or has length >4>4 and in the induced graph $G[V_{2}]$ has tree-width bounded by $b$. While this looks artificial at a first glance, it turns out that there is a constant $b$ such that $(k,4)$-restricted permanent on $(4,b)$-nice graphs describes the natural class $\mathsf{VW}[\mathsf{F}]$. There is a family of $(4,b)$-nice graphs such that the corresponding family of $(k,4)$-restricted permanents is $\mathsf{VW}[\mathsf{F}]$-hard. On the other hand, the $(k,4)$-restricted permanent family is in $\mathsf{VW}[\mathsf{F}]$ for every family of $(4,b)$-nice graphs. Together, this implies:
@@ 1.2 Proof ideas
Proof of Part (1): We prove even a stronger statement for the subexponential version of the linear counting hierarchy. The proof goes via induction on the level of the counting hierarchy. The criteria for some language $B$ being in the $(k+1)$-th level are that there should be some language $A$ in the $k$-th level such that $|\{y\in\{0,1\}^{n}:\langle x,y\rangle\in A\}|>2^{n-1}$. Essentially, for a language $A$ in the $k$-th level, we express $|\{y\in\{0,1\}^{n}:\langle x,y\rangle\in A\}|>2^{n-1}$ as an exponential sum over an algebraic circuit $\chi_{A}(x,y)$, which captures the characteristic function of $A$. Furthermore, one can show that $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is fpt (in an unbounded constant-free setting) iff $\sum_{y}g(\mathbf{X},y)$ has $2^{o(n)}{\rm poly}(m)$ size circuits, where $g$ has a circuit of size $m$; see Theorem˜4.3 and 4.4. Putting these together, one gets that the exponential sum has a subexponential-size constant-free circuit. Lastly, we want to get the information about the highest bit of the sum (which is equivalent to looking at it $\bmod\penalty 10000\ 2^{n}$), which can be efficiently arithmetized. In every step there is polynomial blowup in the size, and hence the size remains subexponential, yielding the desired result. For details, see Theorem˜4.1.
@@
Proof of Part (2): This proof is an adaption of [BÜR09, KOI05] in our context. Take a sequence $(a_{n})_{n}\in\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. We define a multilinear polynomial $A(\mathbf{Y})$ such that the coefficient of $\mathbf{Y}^{\mathbf{j}}$ is the $j$-th bit of $a(n)$, where $\mathbf{j}$ is the binary representation of $j$. Furthermore, checking $a(n,j)=b$ can be done by a subexponential circuit $C(\mathbf{N},\mathbf{J})$, where $\mathbf{N}$ and $\mathbf{J}$ have $\log n$ and $\mathsf{bit}(n)$-many variables capturing $n$ and $j$ respectively. Moreover, one can define $F(\mathbf{N},\mathbf{Y},\mathbf{J})=C(\mathbf{N},\mathbf{J})\cdot\prod_{i}(J_{i}Y_{i}+1-J_{i})$ and show that $A$ can be expressed as an exponential sum over $F(j,\mathbf{N},\mathbf{Y})$! This is clearly a $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ instance, which finally yields that the $\tau$-complexity of $a(n)$ is subpolynomial. A similar proof strategy also holds for the polynomials with coefficients being definable in $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$. For details, see Section˜6.
@@
Proof idea of Theorem˜1.3 Take the Pochhammer polynomial $p_{n}(X)=\prod_{i=1}^{n}(X+i)$. The coefficient of $X^{n-k}$ in $p_{n}$ will be $\sigma_{k}(1,\dots,n)$, where $\sigma_{k}(z_{1},\dots,z_{n})$ is the $k$-th elementary symmetric polynomial in variables $z_{1},\dots,z_{n}$. One can show that $\mathsf{CH}_{\mathsf{lin}}\mathsf{P}$ is closed under polynomially-many additions and multiplications (Theorem˜5.3). Therefore, $(\sigma_{k}(1,\dots,n))_{n\in\mathbb{N},k\leq n}$ is definable in the linear counting hierarchy (see Corollary˜5.4). And by Theorem˜6.1, $(p_{n})_{n\in\mathbb{N}}$ has $n^{o(1)}$-sized constant-free circuits if $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is fixed-parameter tractable. But $p_{n}$ has $n$ distinct integer roots. Assuming the $\tau$-conjecture, $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is not fpt. On the other hand, one can show that when exponential sums over circuits of size $m$ have circuits have size $2^{o(n)}{\rm poly}(m)$, then the $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is fpt, by Theorem 4.4; in other words, $\mathsf{p\text{-}{{\log}\text{-}Expsum}}$ is not fpt implies an exponential lower bound on an exponential sum. This finishes the proof.